# Symmetry Adapted Analysis of Screw Dislocation: Electronic Structure and Carrier Recombination Mechanisms in GaN

### v1

@@ II Symmetry Adapted Basis
For generality we consider a n-fold screw operation that rotates by $2\pi/n$ about the $z$-axis and translates by $mc/n$ along $z$. So a screw symmetry is characterized by the rotation index $m$ and translation index $n$.
@@ III.1 Crystal structure and computational method
Given that the geometry of a screw dislocation inherently breaks translational symmetry, we adopt a quasi-one-dimensional GaN nanowire model. Vacuum layers are applied in the $xy$-plane to isolate the wire, and the lateral surfaces are passivated using pseudo-hydrogen atoms with fractional charges to eliminate dangling bond surface states.[3] To facilitate the symmetry-adapted analysis, the supercell is constructed to enclose complete symmetry orbits, ensuring that for any included atom, all its images generated by the screw operator are also present. The resulting atomic structure is illustrated in Fig. 1.
@@ III.2 Hamiltonian and Band structure
Fig. 3 compares the band structures of ideal and screw-dislocated GaN nanowires. The pristine nanowire exhibits a band gap of 3.68 eV, in good agreement with experimental values (3.51 eV [13, 23]). Conversely, the introduction of a screw dislocation induces six distinct localized states and significantly reduces the band gap to 0.70 eV, consistent with prior findings [3]. Utilizing our symmetry-based analysis, we demonstrate that these localized states originate from the $\mu=0,1,$ and $5$ symmetry channels. Notably, the states flanking the Fermi level (labeled 3 and 4) are exclusively associated with the $\mu=0$ representation.
@@ IV.3 Radiative Recombination Rate
Figure 6: Radiative recombination rate as a function of injected carrier density $n$. The decomposition into different $\Delta\mu$ channels is indicated by dashed lines in different colors for (a) ideal GaN and (b) GaN with a screw dislocation. These comparison shows the existence of screw dislocation significantly suppresses the radiative ability of GaN.
@@
Figure 7: Electrostatic potential distribution and orbital composition analysis of screw-dislocated GaN. (a)-(d) Cross-sectional maps of the electrostatic potential at different $z$-coordinates ($z=0.00,1.30,2.75,$ and $4.13$ Å). (e)-(f) Orbital composition of state 3 and 4: The marker size represents the orbital weight, and colors represent atom types. Electron state is predominantly localized on Nitrogen atoms, while hole state is localized on Gallium atoms.
@@ V Nonradiative recombination
This section distills the essential results for phonon-mediated nonradiative recombination in the screw-dislocated GaN. The derivation follows the Born–Huang expansion and the projection formalism.
@@
The final nonradiative recombination coefficient can be regarded as the superposition of the Huang-Rhys factor ($S_{k}=\omega_{k}\Delta Q_{k}^{2}/2\hbar$) and electron-phonon coupling ($W^{if}$[30]. These two components are contributed by the phonons with different frequencies, which are shown in Fig. 8. It elucidates that while the Huang-Rhys factor involves contributions from phonons across the entire frequency spectrum, the low-frequency regime dominates the electron-phonon coupling. Consequently, the total non-radiative recombination coefficient is essentially determined by the low-frequency phonon modes.
@@
From the phonon-mediated capture calculations we obtain the nonradiative recombination coefficient at energy gap $\Delta E=0.7\,\text{eV}$ and $T=300\,\text{K}$ is:
@@
Our calculated nonradiative recombination coefficient of the dislocation is comparable to the coefficient reported for the $C_{N}$ defect in Ref. [1] ($C_{p}=3.1\times 10^{-9}\,\text{cm}^{3}/\text{s}$), and is consistent with the reduced recombination lifetimes observed in dislocation-rich GaN [3]. These agreements validate our finding that screw dislocations exhibit high nonradiative recombination rates.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China (Grants No. 12188101, No. 12334005, No. 12404089, No. 1227408, No. 124B1003), National Key Research and Development Program of China (Grant No. 2024YFA1409800). Y.X. and H.S. contributed equally to this work.

### v2

@@ II Symmetry Adapted Basis
For generality we consider a n-fold screw operation that rotates by $2\pi/n$ about the $z$-axis and translates by $mc/n$ along $z$. So a screw symmetry is characterized by the rotation index $n$ and translation index $m$.
@@ III.1 Crystal structure and computational method
Given that the geometry of a screw dislocation inherently breaks translational symmetry, we adopt a quasi-one-dimensional GaN nanowire model. Vacuum layers are applied in the $xy$-plane to isolate the wire, and the lateral surfaces are passivated using pseudo-hydrogen atoms with fractional charges to eliminate dangling bond surface states.[8]
[…]
@@ III.2 Hamiltonian and Band structure
Fig. 3 compares the band structures of ideal and screw-dislocated GaN nanowires. The pristine nanowire exhibits a band gap of 3.68 eV, in good agreement with experimental values (3.51 eV [22, 16]). Conversely, the introduction of a screw dislocation induces six distinct localized states and significantly reduces the band gap to 0.70 eV, which is consistent with prior calculations [8] and aligns with the $\sim$0.75 eV deep-level defects experimentally observed in GaN via optical deep-level transient spectroscopy (ODLTS) [23].
The emergence of these discrete states is primarily driven by the localized potential well at the screw dislocation core, which ”traps” electronic states within the bandgap. This strong localization effectively narrows the electronic gap to 0.70 eV. A more detailed physical analysis of how the resulting spatial charge separation further impacts recombination processes is provided in Section IV.4.
Utilizing our symmetry-based analysis, we demonstrate that these localized states originate from the $\mu=0,1,$ and $5$ symmetry channels. Notably, the states flanking the Fermi level (labeled 3 and 4) are exclusively associated with the $\mu=0$ representation.
@@ IV.3 Radiative Recombination Rate
Figure 6: Radiative recombination rate as a function of injected carrier density $n$. The decomposition into different $\Delta\mu$ channels is indicated by dashed lines in different colors for (a) ideal GaN and (b) GaN with a screw dislocation. This comparison shows the existence of screw dislocation significantly suppresses the radiative ability of GaN.
@@
Figure 7: Electrostatic potential distribution and orbital composition analysis of screw-dislocated GaN at the $\Gamma$ point. (a)-(d) Cross-sectional maps of the electrostatic potential at different $z$-coordinates ($z=0.00,1.30,2.75,$ and $4.13$ Å). (e)-(f) Orbital compositions of states 3 and 4: The marker size represents the orbital weight evaluated via Mulliken population analysis, and colors represent atom types. The hole state is predominantly localized on nitrogen atoms, while the electron state is localized on gallium atoms.
@@ V Nonradiative recombination
This section distills the essential results for phonon-mediated nonradiative recombination in the screw-dislocated GaN. Given that the energy gap between band 3 and 4 is the smallest, the multi-phonon non-radiative process is expected to be most efficient and thus primarily takes place between these two dislocation states.
@@
Calculation shows there is a lattice relaxation of the screw dislocation, i.e., the equilibrium structures differ before and after nonradiative recombination. We calculate the partial Huang-Rhys factor ($S_{k}=\omega_{k}\Delta Q_{k}^{2}/2\hbar$[31] at each phonon mode $k$ to reflect the strength of lattice relaxation, as shown in the upper panel of Fig. 8. The total Huang-Rhys factor of 2.6 can be obtained by summing $S_{k}$ at each mode, which is much lower than typical $S$ for carrier capture at point defects in GaN, e.g., $S=10.0$ for hole capture at $C_{\text{N}}^{-}$ [32]. Furthermore, the phonon modes with large contribution to $S_{k}$ contribute minorly to the electron-phonon coupling matrix element (Fig. 8). Since the recombination rate is determined by a joint contribution of lattice relaxation and electron-phonon coupling ($W^{if}$[33], such a mismatch will also reduce the practical contribution to this multiphonon process. As a result, the calculated nonradiative recombination is limited, with a coefficient at energy gap $\Delta E=0.7\,\text{eV}$ and $T=300\,\text{K}$,
@@
Although $S$ of the screw dislocation is smaller than that of $C_{\text{N}}^{-}$, the calculated recombination coefficient is still comparable to that of $C_{\text{N}}^{-}$ in Ref. [32] ($C_{p}=3.1\times 10^{-9}\,\text{cm}^{3}/\text{s}$), which can be ascribed to the smaller energy gap in our nonradiative recombination calculations. The calculated coefficient is consistent with the reduced recombination lifetimes observed in dislocation-rich GaN [8].
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China (Grants No. 12188101, No. 12334005, No. 12404089, No. 1227408, No. 124B1003), National Key Research and Development Program of China (Grant No. 2024YFA1409800). The data that support the findings of this study are available within the article and its Supplemental Material as cited in Ref. [34]. Y.X. and H.S. contributed equally to this work.